# Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication

### v1

@@
Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication
Maysam Behmanesh
LIX, École Polytechnique, IP Paris
&Erkan Turan
LIX, École Polytechnique, IP Paris
Maks Ovsjanikov
LIX, École Polytechnique, IP Paris
Graph alignment—the problem of identifying corresponding nodes across multiple graphs—is fundamental to numerous applications. Most existing unsupervised methods embed node features into latent representations to enable cross-graph comparison without ground-truth correspondences. However, these methods suffer from two critical limitations: the degradation of node distinctiveness due to oversmoothing in GNN-based embeddings, and the misalignment of latent spaces across graphs caused by structural noise, feature heterogeneity, and training instability, ultimately leading to unreliable node correspondences. We propose a novel graph alignment framework that simultaneously enhances node distinctiveness and enforces geometric consistency across latent spaces. Our approach introduces a dual-pass encoder that combines low-pass and high-pass spectral filters to generate embeddings that are both structure-aware and highly discriminative. To address latent space misalignment, we incorporate a geometry-aware functional map module that learns bijective and isometric transformations between graph embeddings, ensuring consistent geometric relationships across different representations. Extensive experiments on graph benchmarks demonstrate that our method consistently outperforms existing unsupervised alignment baselines, exhibiting superior robustness to structural inconsistencies and challenging alignment scenarios. Additionally, comprehensive evaluation on vision-language benchmarks using diverse pretrained models shows that our framework effectively generalizes beyond graph domains, enabling unsupervised alignment of vision and language representations.
@@ 1 Introduction
Graph alignment—also referred to as network alignment or graph matching—is a fundamental problem in machine learning and graph theory, concerned with identifying a correspondence between the nodes of two graphs such that structurally similar or semantically equivalent nodes are matched.
Graph alignment arises in a wide range of application domains, including bioinformatics (e.g., protein interaction networks) (Liao et al., 2009; Singh et al., 2007), social network analysis (Li et al., 2018; Korula and Lattanzi, 2014), computer vision (Liu et al., 2022a; Chen et al., 2025; Wang et al., 2019), and natural language processing (Osman and Barukub, 2020; Guillaume, 2021). Due to its combinatorial nature, graph alignment is computationally challenging, often requiring approximation or heuristic algorithms.
Graph alignment methods are typically classified into three categories based on their alignment strategies: optimization-based, optimal transport–based, and embedding-based approaches. They also vary in the level of supervision required, ranging from unsupervised to semi-supervised—using partial node correspondences—and fully supervised methods. A detailed overview of these categories with related works is provided in Appendix A.
Embedding-based graph alignment methods encode graphs into low-dimensional node representations via Graph Neural Networks (GNNs) (He et al., 2024; Fey et al., 2020; Gao et al., 2021a), followed by alignment through transformations or joint learning with cross-graph regularization. Node matching is then performed using nearest-neighbor search or assignment algorithms, achieving better scalability than optimization-based alternatives.
@@
1) Degradation of node distinctiveness in GNN embeddings. While GNNs generate node representations by aggregating features from local neighborhoods to capture structural information, this aggregation process inherently reduces node distinctiveness. This limitation is particularly problematic for graph alignment, where accurate correspondence identification depends on highly discriminative representations. As embeddings lose uniqueness, alignment becomes increasingly ambiguous and error-prone.
@@
This misalignment arises from multiple sources. First, structural inconsistencies—such as missing or noisy edges—distort neighborhood aggregation during message passing, leading to incompatible embeddings for otherwise corresponding nodes. Second, feature inconsistency across graphs—stemming from differences in user attributes, schema, or data domains—causes graph encoders to embed semantically equivalent nodes into disjoint subspaces. Lastly, GNN-based encoders often exhibit stochasticity in training; different random initializations can yield drastically different embeddings, even on fixed graph inputs (Moschella et al., 2023). Without explicit mechanisms to harmonize or align the latent spaces, these inconsistencies severely hinder cross-graph communication and undermine the reliability of node alignment.
@@
Figure 1: Illustration of limitations in embedding-based graph alignment using synthetic data. $\mathcal{G}_{1}$ is a ring graph with 100 nodes and 2D random features. (a) $\mathcal{G}_{2}$ is identical to $\mathcal{G}_{1}$, yielding well-aligned embeddings. (b) Feature inconsistency introduced through Gaussian noise (std=0.2) causes divergence. (c) Structural inconsistency via 30% edge dropout distorts the embedding alignment. (d) Both graphs are identical, but different training runs produce inconsistent embedding spaces, highlighting instability in unsupervised latent space learning.
In this paper, we introduce GADL—Graph Alignment with Dual-pass encoder and Latent space communication—which builds upon the Graph Autoencoder (GAE) framework (Kipf and Welling, 2016) by incorporating a dual-pass encoding architecture and cross-graph latent communication mechanism tailored for unsupervised graph alignment tasks.
First, to address the degradation of node distinctiveness caused by oversmoothing in GNN neighborhood aggregation, GADL employs a dual-pass GCN encoder that combines low-pass and high-pass spectral filters. The low-pass branch captures structural context, while the high-pass branch preserves fine-grained node distinctiveness. Their concatenation yields embeddings that are both structure-aware and highly discriminative—crucial for accurate graph alignment. Second, to address latent space misalignment across graphs, GADL incorporates a geometry-aware functional map module that learns explicit transformations between different graph embeddings. By enforcing bijectivity and orthogonality constraints, this module ensures that embeddings across graphs become mutually consistent and locally isometric, enabling effective cross-graph communication and alignment without requiring ground-truth node pairs.
Our key contributions are summarized as follows:
@@ 2.1 Graph autoencoder for unsupervised node embedding
Graph autoencoders (GAEs) (Kipf and Welling, 2016) are graph representation learning frameworks used to learn node embeddings in an unsupervised setting. They aim to generate low-dimensional representations that effectively capture both node features and the underlying geometric structure of the graph.
Following the general principle of autoencoders, a GAE consists of two main components: an encoder $q_{\theta}(\mathbf{Z}\mid\mathcal{G})$ that maps the input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$—where $\mathbf{X}\in\mathbb{R}^{|V|\times k}$—into a latent embedding matrix $\mathbf{Z}\in\mathbb{R}^{|V|\times d}$, leveraging both graph structure and node features to learn meaningful representations; and a decoder $p_{\phi}(\mathcal{G}\mid\mathbf{Z})$ that reconstructs the original graph structure and node attributes from these latent embeddings, producing an approximation $\hat{\mathcal{G}}$ of the input graph. The model is trained to minimize a loss function composed of a reconstruction loss $\mathcal{L}_{\text{rec}}$, which measures the difference between $\mathcal{G}$ and $\hat{\mathcal{G}}$, and optionally a regularization term $\mathcal{L}_{\text{reg}}$ on the latent space:
$\mathcal{L}=\mathcal{L}_{\text{rec}}(\mathcal{G},\hat{\mathcal{G}})+\lambda\mathcal{L}_{\text{reg}}(\mathbf{Z}),$
@@ 2.2 Functional map on graphs
Building on the general framework of Deep Geometric Functional Maps (Donati et al., 2020), the functional map formulation is adapted to operate on graph-based latent representations through the following steps:
@@
Given a pair of graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, each domain is associated with a set of descriptor functions, denoted by $\mathcal{F}_{\theta}({\mathcal{G}1})$ and $\mathcal{F}_{\theta}({\mathcal{G}_{2}})$, respectively. A descriptor function is a real-valued function defined on the nodes of a graph—either hand-crafted to capture structural information shared across graphs or learned via neural encoders—producing row feature matrices $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$.
@@ 2. Projection to spectral domain:
For each domain, the spectral basis $\Phi_{*}$ is computed by performing eigendecomposition of the normalized graph Laplacian, defined as $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. The descriptor functions are then projected onto the reduced spectral subspace $\Phi_{*}\in\mathbb{R}^{n_{*}\times r}$, spanned by the first $r$ eigenvectors, resulting in the spectral coefficients $\hat{\mathbf{F}}_{1}=\Phi_{1}^{\top}\mathbf{F}_{1}$, and $\hat{\mathbf{F}}_{2}=\Phi_{2}^{\top}\mathbf{F}_{2}$.
@@
where the second term is the Laplacian commutativity regularizer, which enforces that the functional map $\mathbf{C}_{12}$ approximately commutes with the graph Laplacians, thereby ensuring that $\mathbf{C}_{12}$ preserves their spectral properties.
@@ 3 Method overview
In this section, we introduce the proposed learning-based framework for the graph alignment problem. First, we present an overview of the framework, followed by a discussion of the inherent limitations of learning-based methods in graph alignment problems. Then, we present the motivation and the key innovations in the proposed framework that address these challenges. Figure 2 provides a schematic overview of the proposed framework.
Figure 2: Overview of the proposed framework. Given a pair of input graphs, the model uses a dual-pass encoder with shared parameters to extract node embeddings. To resolve latent space misalignment, a regularized functional map module is introduced to explicitly enforce structural constraints and enable communication between latent spaces. A graph decoder reconstructs the inputs, and the model is optimized with an overall loss. Finally, alignments are estimated via cosine similarity between the embeddings, followed by greedy matching to establish one-to-one correspondences.
@@ 3.1 Overall framework
Given two graphs $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\mathbf{X}_{s})$ and $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$ as input, the objective is to learn meaningful latent representations of the nodes. The proposed architecture consists of two graph autoencoders with shared parameters, each designed to extract latent embeddings $\mathbf{Z}_{s}\in\mathbb{R}^{|V_{s}|\times d}$ and $\mathbf{Z}_{t}\in\mathbb{R}^{|V_{t}|\times d}$ from the respective input graphs. This is achieved by jointly encoding both the graph structure and node attributes through graph encoders $\mathbf{Z}_{s}=f^{s}_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})\in\mathbb{R}^{|V_{s}|\times d}$ and $\mathbf{Z_{t}}=f^{t}_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})\in\mathbb{R}^{|V_{t}|\times d}$, where $\mathbf{A}$ denotes the adjacency matrix encoding the graph structure, and $\mathbf{X}$ contains the node features.
As outlined in the introduction, this framework encounters two fundamental limitations that substantially reduce its effectiveness for graph alignment: loss of node distinctiveness and misaligned latent spaces across graphs in unsupervised settings. In the following sections, we present our proposed framework, which is specifically designed to address these limitations.
@@ 3.2 Graph encoder
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, a graph encoder $\mathbf{Z}=f_{\theta}(\mathbf{X},\mathbf{A})\in\mathbb{R}^{|V|\times d}$ embeds each node $v_{i}\in\mathcal{V}$ into a low-dimensional latent vector $\mathbf{z}_{i}\in\mathbb{R}^{d}$, such that the embeddings of neighboring nodes are encouraged to be similar. While this property allows the encoder to capture the local graph structure effectively, it poses a significant limitation for graph alignment tasks by reducing the distinctiveness of individual nodes—an essential factor for accurately identifying corresponding nodes across graphs.
@@
An ideal node embedding for graph alignment simultaneously achieves two key properties: local consistency, where the embeddings of neighboring nodes are similar such that $\max_{v\in V}\max_{u\in\mathcal{N}(v)}\|h_{v}^{(k)}-h_{u}^{(k)}\|$ is small, and global distinctiveness, where embeddings of distinct nodes are sufficiently different to allow unique identification, i.e., $\min_{v,w\in V}\|h_{v}^{(k)}-h_{w}^{(k)}\|$. As a result, accurate graph alignment requires embedding nodes that balance a fundamental trade-off: they must preserve local similarities to capture the underlying graph structure, while simultaneously preserving node distinctiveness by limiting excessive feature mixing to ensure that each node remains uniquely identifiable.
@@
One of the simple yet effective graph encoders is the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), which extends the convolution operation to graph-structured data. GCNs learn node representations by aggregating information from neighboring nodes, thereby capturing both node fetures and graph structure. A single GCN layer is defined by $\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with added self-loops, $\tilde{\mathbf{D}}$ is the diagonal degree matrix of $\tilde{\mathbf{A}}$, $\mathbf{H}^{(l)}$ is the node feature matrix at the $l$-th layer, $\mathbf{W}^{(l)}$ is the learnable weight matrix of layer $l$, and $\sigma(\cdot)$ is a non-linear activation function such as ReLU.
@@
Spectral interpretation of GCN: In graph signal processing, the graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix and $\mathbf{A}$ is the adjacency matrix. The Laplacian can be decomposed as $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})$ is the matrix of eigenvectors, and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n})$ is the diagonal matrix of eigenvalues. The normalized graph Laplacian is defined as $\mathbf{L}_{\text{sym}}=\mathbf{D}^{-1/2}\mathbf{L}\mathbf{D}^{-1/2}$, whose eigenvalues $\lambda_{i}$ lie within the interval $[0,2]$.
@@
The GCN filter can be expressed as $\tilde{\mathbf{A}}_{\text{GCN},\text{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, with an associated frequency response function $p_{\text{GCN}}(\tilde{\lambda}_{i})=1-\tilde{\lambda}_{i}$. Since the eigenvalues satisfy $\tilde{\lambda}_{i}\in[0,2)$, the response function $p_{\text{GCN}}(\tilde{\lambda}_{i})$ decreases as $\tilde{\lambda}_{i}$ increases, particularly over the range $[0,1]$. This behavior implies that the GCN filter primarily suppresses high-frequency components and thus acts as a low-pass filter in that region. However, for $\tilde{\lambda}_{i}>1$, $p_{\text{GCN}}(\tilde{\lambda}_{i})$ becomes negative, introducing noise and disrupting smoothness. This means GCN is not a completely low-pass filter and can degrade performance due to this issue.
@@
The low-pass graph filter is characterized by $\tilde{\mathbf{A}}_{l,\text{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{l}\tilde{\mathbf{D}}^{-1/2}=\mathbf{I}-\frac{1}{2}\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}\left(\mathbf{I}-\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}$. This formulation reveals that $\tilde{\mathbf{A}}_{l,\text{sym}}$ exhibits a frequency response $p_{l}(\tilde{\lambda}_{i})=1-\frac{1}{2}\tilde{\lambda}_{i}$. The response function is monotonically decreasing over $\tilde{\lambda}_{i}\in[0,2]$, thereby attenuating high-frequency components while preserving smooth graph signals. This design enables the encoder to capture local structural patterns and maintain graph regularity in the learned embeddings. Similarly, the high-pass graph filter is defined by $\tilde{\mathbf{A}}_{h,\text{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{h}\tilde{\mathbf{D}}^{-1/2}=\frac{1}{2}\tilde{\mathbf{L}}_{\text{sym}}=\mathbf{U}\left(\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}.$ This formulation indicates that $\tilde{\mathbf{A}}_{h,\text{sym}}$ acts as a spectral filter with frequency response $p_{h}(\tilde{\lambda}_{i})=\frac{1}{2}\tilde{\lambda}_{i}$. The frequency response $p_{h}(\tilde{\lambda}_{i})$ is a non-negative, monotonically increasing function over the eigenvalue range $\tilde{\lambda}_{i}\in[0,2]$. Consequently, the filter suppresses low-frequency components while amplifying high-frequency components, thereby functioning as a purely high-pass filter (see (Wang et al., 2022a) for details on the frequency response).
@@
Let $\mathbf{z}_{i}^{\text{low}}\in\mathbb{R}^{d_{1}}$ and $\mathbf{z}_{i}^{\text{high}}\in\mathbb{R}^{d_{2}}$ denote the node embeddings of node $i$ in graph $\mathcal{G}$ produced by low-pass and high-pass GCN encoders, respectively and dual-pass embedding is defined as the concatenation $\mathbf{z}_{i}=[\mathbf{z}_{i}^{\text{low}}\,\|\,\mathbf{z}_{i}^{\text{high}}]\in\mathbb{R}^{d_{1}+d_{2}}$. Assume both input graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ use this encoder architecture to compute embeddings $\mathbf{z}_{1}$ and $\mathbf{z}_{2}$, respectively. Then, the dual-pass GCN encoder constitutes an ideal node embedding for graph alignment by satisfying the following properties:
@@
    Spectral locality preservation: the embedding $\mathbf{z}_{i}$ preserves neighborhood similarity comparably to $\mathbf{z}_{i}^{\text{low}}$.

### v2

@@
Abstract
Graph alignment, the problem of identifying corresponding nodes across multiple graphs, is fundamental to numerous applications. Most existing unsupervised methods embed node features into latent representations to enable cross-graph comparison without ground-truth correspondences. However, these methods suffer from two critical limitations: the degradation of node distinctiveness due to oversmoothing in GNN-based embeddings, and the misalignment of latent spaces across graphs caused by structural noise, feature heterogeneity, and training instability, ultimately leading to unreliable node correspondences. We propose a novel graph alignment framework that simultaneously enhances node distinctiveness and enforces geometric consistency across latent spaces. Our approach introduces a dual-pass encoder that combines low-pass and high-pass spectral filters to generate embeddings that are both structure-aware and highly discriminative. To address latent space misalignment, we incorporate a geometry-aware functional map module that learns bijective and isometric transformations between graph embeddings, ensuring consistent geometric relationships across different representations. Extensive experiments on graph benchmarks demonstrate that our method consistently outperforms existing unsupervised alignment baselines, exhibiting superior robustness to structural inconsistencies and challenging alignment scenarios. Additionally, comprehensive evaluation on vision-language benchmarks using diverse pretrained models shows that our framework effectively generalizes beyond graph domains, enabling unsupervised alignment of vision and language representations.
@@ 1 Introduction
Graph alignment, also referred to as network alignment or graph matching, is a fundamental problem in machine learning and graph theory, concerned with identifying a correspondence between the nodes of two graphs such that structurally similar or semantically equivalent nodes are matched.
Graph alignment arises in a wide range of application domains, including bioinformatics (e.g., protein interaction networks) (Liao et al., 2009; Singh et al., 2007), social network analysis (Li et al., 2018; Korula & Lattanzi, 2014), computer vision (Liu et al., 2022a; Chen et al., 2025; Wang et al., 2019), and natural language processing (Osman & Barukub, 2020; Guillaume, 2021). Due to its combinatorial nature, graph alignment is computationally challenging, often requiring approximation or heuristic algorithms.
Graph alignment methods are typically classified into three categories based on their alignment strategies: optimization-based, optimal transport–based, and embedding-based approaches. They also vary in the level of supervision required, ranging from unsupervised to semi-supervised, using partial node correspondences, and fully supervised methods. A detailed overview of these categories with related works is provided in Appendix A.
Embedding-based graph alignment methods encode graphs into low-dimensional node representations via Graph Neural Networks (GNNs) (He et al., 2024; Fey et al., 2020; Gao et al., 2021b), followed by alignment through transformations or joint learning with cross-graph regularization. Node matching is then performed using nearest-neighbor search or assignment algorithms, achieving better scalability than optimization-based alternatives.
@@
1) Degradation of node distinctiveness in GNN embeddings. While GNNs capture structural information by aggregating neighborhood features, this process inherently reduces node distinctiveness. This limitation is particularly problematic for graph alignment, where accurate correspondence identification depends on highly discriminative representations. As embeddings lose uniqueness, alignment becomes increasingly ambiguous and error-prone.
@@
This misalignment arises from multiple sources. First, structural inconsistencies, such as missing or noisy edges, distort neighborhood aggregation during message passing, leading to incompatible embeddings for otherwise corresponding nodes. Second, feature inconsistency across graphs, stemming from differences in user attributes, schema, or data domains, causes graph encoders to embed semantically equivalent nodes into disjoint subspaces. Lastly, GNN-based encoders often exhibit stochasticity in training; different random initializations can yield drastically different embeddings, even on fixed graph inputs (Moschella et al., 2023). Without explicit mechanisms to harmonize or align the latent spaces, these inconsistencies severely hinder cross-graph communication and undermine the reliability of node alignment.
@@
Figure 1: Limitations of embedding-based graph alignment on synthetic data. $\mathcal{G}_{1}$ is a ring graph with 100 nodes and 2D random features. (a) $\mathcal{G}_{2}$ is identical to $\mathcal{G}_{1}$, yielding well-aligned embeddings. (b) Feature inconsistency introduced through Gaussian noise (std=0.2) causes divergence. (c) Structural inconsistency via 30% edge dropout distorts the embedding alignment. (d) Identical graphs with different training runs show embedding instability, highlighting unsupervised learning limitations.
In this paper, we introduce GADL, Graph Alignment with Dual-pass encoder and Latent space communication, which builds upon the Graph Autoencoder (GAE) framework (Kipf & Welling, 2016) by incorporating a dual-pass encoding architecture and cross-graph latent communication mechanism tailored for unsupervised graph alignment tasks.
First, to address the degradation of node distinctiveness caused by oversmoothing in GNN neighborhood aggregation, GADL employs a dual-pass GCN encoder that combines low-pass and high-pass spectral filters. The low-pass branch captures structural context, while the high-pass branch preserves fine-grained node distinctiveness. Their concatenation yields embeddings that are both structure-aware and highly discriminative, crucial for accurate graph alignment. Second, to address latent space misalignment across graphs, GADL incorporates a geometry-aware functional map module that learns explicit transformations between different graph embeddings. By enforcing bijectivity and orthogonality constraints, this module ensures that embeddings across graphs become mutually consistent and locally isometric, enabling effective cross-graph communication and alignment without requiring ground-truth node pairs. Our key contributions can be summarized as:
@@ 2.1 Graph autoencoder for unsupervised node embedding
Graph autoencoders (GAEs) (Kipf & Welling, 2016) learn node embeddings in an unsupervised setting, generating low-dimensional representations that capture both node features and graph structure. Following the general principle of autoencoders, a GAE consists of two main components: an encoder $q_{\theta}(\mathbf{Z}\mid\mathcal{G})$ that maps the input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, where $\mathbf{X}\in\mathbb{R}^{|V|\times k}$, into a latent embedding matrix $\mathbf{Z}\in\mathbb{R}^{|V|\times d}$, leveraging both graph structure and node features to learn meaningful representations; and a decoder $p_{\phi}(\mathcal{G}\mid\mathbf{Z})$ that reconstructs the original graph structure and node attributes from these latent embeddings, producing an approximation $\hat{\mathcal{G}}$ of the input graph. The model is trained to minimize a loss function composed of a reconstruction loss $\mathcal{L}_{\mathrm{rec}}$, which measures the difference between $\mathcal{G}$ and $\hat{\mathcal{G}}$, and optionally a regularization term $\mathcal{L}_{\mathrm{reg}}$ on the latent space:
$\mathcal{L}=\mathcal{L}_{\mathrm{rec}}(\mathcal{G},\hat{\mathcal{G}})+\lambda\mathcal{L}_{\mathrm{reg}}(\mathbf{Z}),$ (1)
@@ 2.2 Functional map on graphs
Building on the general framework of Deep Geometric Functional Maps (Donati et al., 2020), the functional map formulation is adapted to operate on graph-based latent representations through:
@@
Given a pair of graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, each is associated with a set of descriptor functions, denoted by $\mathcal{F}_{\theta}({\mathcal{G}1})$ and $\mathcal{F}_{\theta}({\mathcal{G}_{2}})$, respectively. A descriptor function is a real-valued function defined on the nodes of a graph, either hand-crafted to capture structural information shared across graphs or learned via neural encoders, producing row feature matrices $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$.
@@ 2. Projection to spectral domain:
For each domain, the spectral basis $\Phi_{*}$ is computed via eigendecomposition of the normalized graph Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$. Descriptor functions are then projected onto the reduced spectral subspace $\Phi_{*}\in\mathbb{R}^{n_{*}\times r}$, spanned by the first $r$ eigenvectors, resulting in the spectral coefficients $\hat{\mathbf{F}}_{1}=\Phi_{1}^{\top}\mathbf{F}_{1}$, and $\hat{\mathbf{F}}_{2}=\Phi_{2}^{\top}\mathbf{F}_{2}$.
@@
where the second term is the Laplacian commutativity regularizer, enforcing that $\mathbf{C}_{12}$ approximately commutes with the graph Laplacians to preserve spectral properties.
@@ 3 Method overview
As established in the introduction, learning-based frameworks for graph alignment suffer from two fundamental limitations that significantly impair their performance: loss of node distinctiveness through feature aggregation and misaligned latent spaces in unsupervised cross-graph scenarios. In the following, we present the proposed framework that addresses these challenges through architectural innovations that preserve node distinguishability while enforcing embedding space alignment.
Figure 2: Overview of the proposed framework. Given input graphs, the model uses a dual-pass encoder with shared parameters to extract node embeddings. A regularized functional map module resolves latent space misalignment by enforcing structural constraints and enabling cross-space communication. A graph decoder reconstructs the inputs, and the model is optimized with an overall loss. Finally, alignments are estimated via cosine similarity and greedy matching.
@@ 3.1 Overall framework
Given two graphs $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\mathbf{X}_{s})$ and $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$, the framework employs a dual-pass encoder with shared parameters $\theta$ to extract meaningful node representations. The encoder processes both graphs simultaneously, generating latent embeddings $\mathbf{Z}_{s}=f_{\theta}(\mathbf{X}_{s},\mathbf{A}_{s})\in\mathbb{R}^{|V_{s}|\times d}$ and $\mathbf{Z}_{t}=f_{\theta}(\mathbf{X}_{t},\mathbf{A}_{t})\in\mathbb{R}^{|V_{t}|\times d}$ by jointly encoding graph structure and node attributes. To address latent space misalignment, a regularized functional map module enforces structural constraints and enables communication between the embedding spaces. Figure 2 provides a schematic overview.
@@ 3.2 Graph encoder
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$, a graph encoder $\mathbf{Z}=f_{\theta}(\mathbf{X},\mathbf{A})\in\mathbb{R}^{|V|\times d}$ embeds each node $v_{i}\in\mathcal{V}$ into a latent vector $\mathbf{z}_{i}\in\mathbb{R}^{d}$, such that the embeddings of neighboring nodes are encouraged to be similar. While this property allows the encoder to capture the local graph structure effectively, it poses a significant limitation for graph alignment tasks by reducing the distinctiveness of individual nodes, an essential factor for accurately identifying corresponding nodes across graphs.
@@
An ideal node embedding for graph alignment achieves two properties: local consistency, where neighboring node embeddings are similar ($\max_{v\in V}\max_{u\in\mathcal{N}(v)}\|h_{v}^{(k)}-h_{u}^{(k)}\|$ is small), and global distinctiveness, where distinct nodes have sufficiently different embeddings ($\min_{v,w\in V}\|h_{v}^{(k)}-h_{w}^{(k)}\|$ is large). Graph alignment thus requires embedding nodes that balance a fundamental trade-off: preserving local similarities to capture structure while maintaining node distinctiveness for unique identification.
@@
One of the simple yet effective graph encoders is the Graph Convolutional Network (GCN) (Kipf & Welling, 2017), which extends convolution to graph-structured data by aggregating neighboring node information to capture both features and structure. A single GCN layer is defined by $\mathbf{H}^{(l+1)}=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}\right)$, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops, $\tilde{\mathbf{D}}$ is the degree matrix, $\mathbf{H}^{(l)}$ and $\mathbf{W}^{(l)}$ are the feature and weight matrices, and $\sigma(\cdot)$ is the activation function.
@@
Spectral interpretation of GCN: In graph signal processing, the graph Laplacian is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix and $\mathbf{A}$ is the adjacency matrix. The Laplacian can be decomposed as $\mathbf{L}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, where $\mathbf{U}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})$ is the matrix of eigenvectors, and $\mathbf{\Lambda}=\operatorname{diag}(\lambda_{1},\ldots,\lambda_{n})$ is the diagonal matrix of eigenvalues. The normalized graph Laplacian is defined as $\mathbf{L}_{\mathrm{sym}}=\mathbf{D}^{-1/2}\mathbf{L}\mathbf{D}^{-1/2}$, whose eigenvalues $\lambda_{i}$ lie within the interval $[0,2]$.
@@
The GCN filter can be expressed as $\tilde{\mathbf{A}}_{\mathrm{GCN},\mathrm{sym}}=\mathbf{I}-\tilde{\mathbf{L}}_{\mathrm{sym}}=\mathbf{U}(\mathbf{I}-\tilde{\mathbf{\Lambda}})\mathbf{U}^{\top}$, with an associated frequency response function $p_{\mathrm{GCN}}(\tilde{\lambda}_{i})=1-\tilde{\lambda}_{i}$. Since the eigenvalues satisfy $\tilde{\lambda}_{i}\in[0,2)$, the response function $p_{\mathrm{GCN}}(\tilde{\lambda}_{i})$ decreases as $\tilde{\lambda}_{i}$ increases, particularly over the range $[0,1]$. This behavior implies that the GCN filter primarily suppresses high-frequency components and thus acts as a low-pass filter in that region. However, for $\tilde{\lambda}_{i}>1$, $p_{\mathrm{GCN}}(\tilde{\lambda}_{i})$ becomes negative, introducing noise and disrupting smoothness. This means GCN is not a completely low-pass filter and can degrade performance due to this issue.
@@
The low-pass graph filter is characterized by $\tilde{\mathbf{A}}_{l,\mathrm{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{l}\tilde{\mathbf{D}}^{-1/2}=\mathbf{I}-\frac{1}{2}\tilde{\mathbf{L}}_{\mathrm{sym}}=\mathbf{U}\left(\mathbf{I}-\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}$. This formulation reveals that $\tilde{\mathbf{A}}_{l,\mathrm{sym}}$ exhibits a frequency response $p_{l}(\tilde{\lambda}_{i})=1-\frac{1}{2}\tilde{\lambda}_{i}$. The response function is monotonically decreasing over $\tilde{\lambda}_{i}\in[0,2]$, thereby attenuating high-frequency components while preserving smooth graph signals. This enables capturing local structural patterns and maintaining graph regularity in embeddings. Similarly, the high-pass graph filter is defined by $\tilde{\mathbf{A}}_{h,\mathrm{sym}}=\tilde{\mathbf{D}}^{-1/2}\tilde{\mathbf{A}}_{h}\tilde{\mathbf{D}}^{-1/2}=\frac{1}{2}\tilde{\mathbf{L}}_{\mathrm{sym}}=\mathbf{U}\left(\frac{1}{2}\tilde{\mathbf{\Lambda}}\right)\mathbf{U}^{\top}.$ This formulation indicates that $\tilde{\mathbf{A}}_{h,\mathrm{sym}}$ acts as a spectral filter with frequency response $p_{h}(\tilde{\lambda}_{i})=\frac{1}{2}\tilde{\lambda}_{i}$. The response $p_{h}(\tilde{\lambda}_{i})$ monotonically increases over $\tilde{\lambda}_{i}\in[0,2]$, suppressing low frequencies while amplifying high frequencies, functioning as a high-pass filter (see (Wang et al., 2022a)).
@@
Let $\mathbf{z}_{i}^{\mathrm{low}}\in\mathbb{R}^{d_{1}}$ and $\mathbf{z}_{i}^{\mathrm{high}}\in\mathbb{R}^{d_{2}}$ denote node embeddings from low-pass and high-pass GCN encoders, respectively, and dual-pass embedding is defined as the concatenation $\mathbf{z}_{i}=[\mathbf{z}_{i}^{\mathrm{low}}\,\|\,\mathbf{z}_{i}^{\mathrm{high}}]\in\mathbb{R}^{d_{1}+d_{2}}$. Using this architecture for both graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, the dual-pass GCN encoder provides ideal node embeddings for graph alignment by satisfying:
@@
    Spectral locality preservation: the embedding $\mathbf{z}_{i}$ preserves neighborhood similarity comparably to $\mathbf{z}_{i}^{\mathrm{low}}$.